- Refurbishment of the power outlets ongoing (EN-EL / James Devine). This will last the whole week. Precautions will be taken no to cut the power for the WISARD setup (ongoing tests this and next week).
- The overhead crane remote is back in the cuboard. PR214 overhead crane can be used by those who have the license again.
- Re-zoning of the overhead crane PR214 in the WISARD area by the transport team (Yann Seraphin) and dewar handling to WISARD on Tuesday 27 Jan.
- EPC platform:
- Construction work of the EPC platform (508 side) ongoing. Crane blockage foreseen next week 4-5(-6) Feb for the dismounting of the scaffolding.
- Helmets have been moved to the left after entrance to the 1st floor of 197. Final position will be on the new cage of the EPC platform once the installation has finished.
- Restricted access from the entrance to the hall from the 508 side during the EPC platform construction work. Visits to the hall from the SAS side (helmets available). Estimate until 6 Feb.
- Request Rustan Calming to store 16 racks for the EPC platform already to the hall: Asked to wait until installation starts (no space in the hall). (Erwin)
- Dust from the construction work is provoking false fire alarms: the fire brigade asked to keep the doors to 197 and the old Control Room closed. Signs have been put in place. Reminder: The key for the door towards 197-1 power room is SC40. Nr 33 in the ISOLDE TRAKA cupboard.
- Cable activities in the hall:
- New power cables distribution for the CHROME RP detectors. New distribution box on the 170 hall wall at the Salve side next to the RP buffer zone. Cables to the HEBT lines and tunnel side (routing through the HIE mezzanine). Partial cabling to the RILIS zone (final installation after RILIS work finishes).
- Power cable for warming gas bottles (new flamable gas rack outside 170, Rte Einstein side) foreseen in parallel.
Contact: Jerome Cantet EN-EL-SW
- Recuperation bins for metal, etc. presently in front of B198 (Cryo) to be moved in front of the (to be) diesel position at B199. With the restricted and difficult access to the IBDRS side this is a more logic and useful position. Thierry Dupont (Cryo TSO 198/199) has been informed and is OK (Hanne).
- 508 access road:
- Installation of concrete slab for MULTIPAC - need to close access road for 4 days - VIC tomorrow 28/1, asked to do this Thursday 29/1 and 30/1 -> I will give feedback based on this meeting
- remark Erwin: Would this not clash with the recuperation of the old ventilation unit material bins next to 197 as well as with the construction work on the EPC platform (contractors van / material)?
